# Learning-Augmented Robust Algorithmic Recourse

### v2

@@ C.1 Non-convexity of Robust Recourse for Linear Models
    To conclude with a contradiction, it suffices to show that $|\theta^{\prime}[j]|>|\theta^{\prime}[i]|$, as this would imply $x^{\prime}\cdot\theta^{\prime}>x^{*}\cdot\theta^{\prime}$, contradicting the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$, since $x^{\prime}$ would require the same cost as $x^{*}$ but it would yield a greater inner product. We consider three possible scenarios: i) If Algorithm 4 in line 12 chose to change dimension $i$ facing adversary $\theta^{\prime}[i]$ before considering dimension $j$ and adversary $\theta^{\prime}[j]$, then the fact that $|x^{*}[i]-x_{0}[i]|>|x_{r}[i]-x_{0}[i]|$ implies that the algorithm did not change coordinate $i$ as much as $x^{*}$ and it must have terminated after that via line 16; this would suggest that dimension $j$ and adversary $\theta[j]$ would never be reached after that, contradicting the fact that $|x^{*}[j]-x_{0}[j]|<|x_{r}[j]-x_{0}[j]|$. ii) If Algorithm 4 in line 12 chose to change dimension $j$ facing adversary $\theta^{\prime}[j]$ and later on also considered dimension $i$ and adversary $\theta^{\prime}[i]$, then Lemma C.4 suggests that $|\theta^{\prime}[j]|>|\theta^{\prime}[i]|$, once again leading to a contradiction. Finally, iii) if Algorithm 4 in line 12 chose to change dimension $j$ facing adversary $\theta^{\prime}[j]$ and never ended up considering dimension $i$ even though $|\theta^{\prime}[j]|<|\theta^{\prime}[i]|$, this suggests that $i$ was removed from the Active set during the execution of the algorithm, which implies that $x_{r}[i]=0$ and $|\theta_{0}[i]<\alpha$, so moving further away from $x_{0}[i]$ would actually hurt the inner product because the adversary can flip the sign of $\theta^{\prime}[i]$ via a change of $\alpha$. The fact that $x^{*}$ actually moved dimension $i$ further away then again contradicts the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$.
@@
  • Case 2: $\|x^{*}-x_{0}\|_{1}<\|x_{r}-x_{0}\|_{1}$. In this case, we can infer that for some $i$ we have $|x^{*}[i]-x_{0}[i]|<|x_{r}[i]-x_{0}[i]|$, i.e., $x^{*}$ determined that the increase of the inner product achieved by moving $x^{*}[i]$ further away from $x_{0}[i]$ and closer to $x_{r}[i]$ was not worth the cost suffered by this increase. However, note that as we discussed in Observation C.1, $J(\cdot)$ is a decreasing function of the inner product. Also note that, since Algorithm 4 changes a coordinate of the recourse only if it increases the inner product, there must be some point in time during the execution of the algorithm when the inner product of $x^{\prime}\cdot\theta^{\prime}$ was at least as high as the inner product of $x^{*}$ with the adversarial response to $x^{*}$. Nevertheless, line 13 determined that this change would decrease the objective value $J(\cdot)$. If we specifically consider the last dimension $j$ changed by the algorithm, using Lemma C.4, we can infer that the value of $|\theta^{\prime}[j]|$ at the time of this change was less than the value of $|\theta^{\prime}[i]|$ for the dimension $i$ satisfying $|x^{*}[i]-x_{0}[i]|<|x_{r}[i]-x_{0}[i]|$; this is due to the fact that the algorithm chose to change $i$ weakly earlier than $j$. As a result, since line 13 determined that the increase of cost was outweighed by the increase in the inner product even though $|\theta^{\prime}[j]|\leq|\theta^{\prime}[i]|$, the inner product is greater, and $J(\cdot)$ is convex in the latter, this implies that increasing the value of $x^{*}[i]$ would also decrease the objective, thus leading to a contradiction of the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$.
@@
Appendix D Omitted Details from Section 4
In this section, we provide additional results and analysis that were omitted from Section 4 due to space constraints. In Section D.1, we provide additional details on the running time of our algorithm as well as what values were chosen for some of the hyperparameters. Section D.2 provides error bars for the robustness consistency tradeoffs generated by our algorithm. Section D.3 details how parameter changes affect our results. Section D.4 studies the effect of post-processing and actionability of the generated recourse on performance. In Section D.5, we provide baselines for our smoothness analysis in Section 4.1. Section D.6 studies the trade-off between cost and validity when there is uncertainty in the value of $\alpha$. Finally, Section D.7 provides additional results for a larger dataset.
@@ Appendix D Omitted Details from Section 4
Table 1: $\lambda$ that maximize the validity with respect to the original model $\theta_{0}$ for each dataset. The other choices of parameters are mentioned in Section D.1.
D.1 Additional Experimental Details
The experiments were conducted on two laptops: an Apple M1 Pro and a 2.2 GHz 6-Core Intel Core i7. Average runtime (across datasets/model) to generate a robust recourse for Algorithm 1, ROAR [70], and RBR [56] are 0.001, 1,0, and 40 seconds, respectively. The total runtime of Algorithm 1 to generate Figure 1 was 45-60 minutes, depending on the subfigure.
@@ D.1 Additional Experimental Details
In our experiment on smoothness in Section 4.1, we created the future model using a modified dataset similar to [70]. To produce the altered synthetic data, we employed the same method outlined in Section 4, but we changed the mean of the Gaussian distribution for class 0. The new distribution is $x\sim N(\mu^{\prime}_{0},\Sigma_{y})$, where $\mu^{\prime}_{0}$ is equal to $\mu_{0}$ + $\alpha,0]^{T}$, while $\mu^{\prime}_{1}$ remained unchanged at $\mu_{1}$. We used this new distribution to learn a model for the correct prediction. The German credit dataset [34] is available in two versions, with the second one [27] fixing coding errors found in the first. This dataset exemplifies a shift due to data correction. We used the second dataset to learn the model for the correct prediction. The Small Business Administration dataset [52], which contains data on 2,102 small business loans approved in California from 1989 to 2012, demonstrates temporal shifts. We split this dataset into two parts: data points before 2006 form the original dataset, while those from 2006 onwards constitute the shifted dataset. We used the shifted dataset to learn a model for the correct prediction.
In our cost versus worst-case validity experiments in Section 4.2, motivated by reasons for data shift and given access to different versions of the datasets, we follow prior work [70, 56] and compute a model on a shifted version of the dataset and measure validity against this new model. For the synthetic dataset, the shifted data is achieved by changing the mean and variance in the data distribution. For the German Credit dataset, the data shift is due to data correction. For the Small Business Administration dataset, the shift is temporal. See [70] for more details.
D.2 Error Bars for Robustness-Consistency Trade-off
@@ D.3 Effect of the Parameters
D.4 Actionable Features and Post-processing
Similar to prior work [70, 56], our algorithm does not directly handle categorical features, and the recourse provided by our algorithm can violate potential constraints or restrictions that might exist on some features. In this section, we study how the performance of our algorithm can be affected under these situations.
In particular, in this section, we repeat the experiments on robustness-consistency trade-off from Section 4.1 but post-process the recourses generated by Algorithm 1 to satisfy categorical and actionable features. In particular, we first remove all non-actionable or immutable features [28] from the dataset, compute the recourses using Algorithm 1, and then project the recourse to the feasible region to satisfy the constraints posed by categorical features.
@@ D.4 Actionable Features and Post-processing
D.5 Baselines for Smoothness Results
In this section, we revisit the smoothness results in Section 4.1 by including a baseline. For any $\beta$, the optimization problem in Equation 7 can be solved by the type of gradient-based techniques that are common in adversarial training [51]. Since ROAR is one such algorithm, we can modify it by changing its objective function and use it as a baseline to Algorithm 1.
@@ D.5 Baselines for Smoothness Results
D.6 Computing Robust Recourse
@@
In this section, we repeat the experiments in Section 4.2 but also include comparisons with RBR [56]. Figure 10 depicts the trade-off between worst-case validity and cost for all datasets (rows) and models (columns). Since RBR does not have the same parameters as our algorithm and ROAR, the trade-off for RBR is obtained by replicating their experiments and setting the ambiguity sizes to $\epsilon_{0},\epsilon_{1}\in[0,1]$ with increments of 0.5, and the maximum recourse cost $\delta=\|x_{0}-x_{r}\|_{1}+\delta_{+}$ to $\delta_{+}\in[0,1]$ with increments of 0.2.
We observe that the trade-offs for RBR are Pareto dominated by our algorithm, except for the synthetic dataset, where RBR achieves a high validity with a smaller cost. In addition, unlike our algorithm, RBR sometimes achieves very low validity, even lower than ROAR.
D.7 Experiments on Larger Datasets
@@
We use the ACSIncome-CA [19] dataset for experiments in this section. This dataset originally consisted of 195,665 data points and 10 features, 7 of which are categorical and have been one-hot encoded. However, to lower the runtime, we subsampled the dataset to include 50,000 data points and removed the categorical feature "occupation (OCCP)", as it contains more than 500 different occupations. This resulted in more than 250 features after one-hot encoding.

### v3

@@ B.1 Non-convexity of Robust Recourse for Linear Models
    To conclude with a contradiction, it suffices to show that $|\theta^{\prime}[j]|>|\theta^{\prime}[i]|$, as this would imply $x^{\prime}\cdot\theta^{\prime}>x^{*}\cdot\theta^{\prime}$, contradicting the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$, since $x^{\prime}$ would require the same cost as $x^{*}$ but it would yield a greater inner product. We consider three possible scenarios: i) If Algorithm 4 in line 12 chose to change dimension $i$ facing adversary $\theta^{\prime}[i]$ before considering dimension $j$ and adversary $\theta^{\prime}[j]$, then the fact that $|x^{*}[i]-x_{0}[i]|>|x_{r}[i]-x_{0}[i]|$ implies that the algorithm did not change coordinate $i$ as much as $x^{*}$ and it must have terminated after that via line 16; this would suggest that dimension $j$ and adversary $\theta[j]$ would never be reached after that, contradicting the fact that $|x^{*}[j]-x_{0}[j]|<|x_{r}[j]-x_{0}[j]|$. ii) If Algorithm 4 in line 12 chose to change dimension $j$ facing adversary $\theta^{\prime}[j]$ and later on also considered dimension $i$ and adversary $\theta^{\prime}[i]$, then Lemma B.5 suggests that $|\theta^{\prime}[j]|>|\theta^{\prime}[i]|$, once again leading to a contradiction. Finally, iii) if Algorithm 4 in line 12 chose to change dimension $j$ facing adversary $\theta^{\prime}[j]$ and never ended up considering dimension $i$ even though $|\theta^{\prime}[j]|<|\theta^{\prime}[i]|$, this suggests that $i$ was removed from the Active set during the execution of the algorithm, which implies that $x_{r}[i]=0$ and $|\theta_{0}[i]<\alpha$, so moving further away from $x_{0}[i]$ would actually hurt the inner product because the adversary can flip the sign of $\theta^{\prime}[i]$ via a change of $\alpha$. The fact that $x^{*}$ actually moved dimension $i$ further away then again contradicts the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$.
@@
  • Case 2: $\|x^{*}-x_{0}\|_{1}<\|x_{r}-x_{0}\|_{1}$. In this case, we can infer that for some $i$ we have $|x^{*}[i]-x_{0}[i]|<|x_{r}[i]-x_{0}[i]|$, i.e., $x^{*}$ determined that the increase of the inner product achieved by moving $x^{*}[i]$ further away from $x_{0}[i]$ and closer to $x_{r}[i]$ was not worth the cost suffered by this increase. However, note that as we discussed in Observation B.2, $J(\cdot)$ is a decreasing function of the inner product. Also note that, since Algorithm 4 changes a coordinate of the recourse only if it increases the inner product, there must be some point in time during the execution of the algorithm when the inner product of $x^{\prime}\cdot\theta^{\prime}$ was at least as high as the inner product of $x^{*}$ with the adversarial response to $x^{*}$. Nevertheless, line 13 determined that this change would decrease the objective value $J(\cdot)$. If we specifically consider the last dimension $j$ changed by the algorithm, using Lemma B.5, we can infer that the value of $|\theta^{\prime}[j]|$ at the time of this change was less than the value of $|\theta^{\prime}[i]|$ for the dimension $i$ satisfying $|x^{*}[i]-x_{0}[i]|<|x_{r}[i]-x_{0}[i]|$; this is due to the fact that the algorithm chose to change $i$ weakly earlier than $j$. As a result, since line 13 determined that the increase of cost was outweighed by the increase in the inner product even though $|\theta^{\prime}[j]|\leq|\theta^{\prime}[i]|$, the inner product is greater, and $J(\cdot)$ is convex in the latter, this implies that increasing the value of $x^{*}[i]$ would also decrease the objective, thus leading to a contradiction of the fact that $x^{*}\in\arg\min_{x^{\prime}\in\mathcal{X}}\penalty 10000\ \max_{\theta^{\prime}\in\Theta_{\alpha}}J(x^{\prime},\theta^{\prime})$.
@@
Appendix C Omitted Details from Section 4
In this section, we provide additional results and analysis that were omitted from Section 4 due to space constraints. In Section C.1, we provide additional details on the running time of our algorithm as well as what values were chosen for some of the hyperparameters. Section C.2 provides error bars for the robustness consistency tradeoffs generated by our algorithm. Section C.3 details how parameter changes affect our results. Section C.4 studies the effect of post-processing and actionability of the generated recourse on performance. In Section C.5, we provide baselines for our smoothness analysis in Section 4.1. Section C.6 studies the trade-off between cost and validity when there is uncertainty in the value of $\alpha$. Finally, Section C.7 provides additional results for a larger dataset.
@@ Appendix C Omitted Details from Section 4
Table 1: $\lambda$ that maximize the validity with respect to the original model $\theta_{0}$ for each dataset. The other choices of parameters are mentioned in Section C.1.
C.1 Additional Experimental Details
The experiments were conducted on two laptops: an Apple M1 Pro and a 2.2 GHz 6-Core Intel Core i7. Average runtime (across datasets/model) to generate a robust recourse for Algorithm 1, ROAR [Upadhyay et al., 2021], and RBR [Nguyen et al., 2022] are 0.001, 1,0, and 40 seconds, respectively. The total runtime of Algorithm 1 to generate Figure 1 was 45-60 minutes, depending on the subfigure.
@@ C.1 Additional Experimental Details
In our experiment on smoothness in Section 4.1, we created the future model using a modified dataset similar to [Upadhyay et al., 2021]. To produce the altered synthetic data, we employed the same method outlined in Section 4, but we changed the mean of the Gaussian distribution for class 0. The new distribution is $x\sim N(\mu^{\prime}_{0},\Sigma_{y})$, where $\mu^{\prime}_{0}$ is equal to $\mu_{0}$ + $\alpha,0]^{T}$, while $\mu^{\prime}_{1}$ remained unchanged at $\mu_{1}$. We used this new distribution to learn a model for the correct prediction. The German credit dataset [Hofmann,Hans, 1994] is available in two versions, with the second one [Grömping, 2020] fixing coding errors found in the first. This dataset exemplifies a shift due to data correction. We used the second dataset to learn the model for the correct prediction. The Small Business Administration dataset [Min Li and Taylor, 2018], which contains data on 2,102 small business loans approved in California from 1989 to 2012, demonstrates temporal shifts. We split this dataset into two parts: data points before 2006 form the original dataset, while those from 2006 onwards constitute the shifted dataset. We used the shifted dataset to learn a model for the correct prediction.
In our cost versus worst-case validity experiments in Section 4.2, motivated by reasons for data shift and given access to different versions of the datasets, we follow prior work [Upadhyay et al., 2021, Nguyen et al., 2022] and compute a model on a shifted version of the dataset and measure validity against this new model. For the synthetic dataset, the shifted data is achieved by changing the mean and variance in the data distribution. For the German Credit dataset, the data shift is due to data correction. For the Small Business Administration dataset, the shift is temporal. See [Upadhyay et al., 2021] for more details.
C.2 Error Bars for Robustness-Consistency Trade-off
@@ C.3 Effect of the Parameters
C.4 Actionable Features and Post-processing
Similar to prior work [Upadhyay et al., 2021, Nguyen et al., 2022], our algorithm does not directly handle categorical features, and the recourse provided by our algorithm can violate potential constraints or restrictions that might exist on some features. In this section, we study how the performance of our algorithm can be affected under these situations.
In particular, in this section, we repeat the experiments on robustness-consistency trade-off from Section 4.1 but post-process the recourses generated by Algorithm 1 to satisfy categorical and actionable features. In particular, we first remove all non-actionable or immutable features [Guldogan et al., 2023] from the dataset, compute the recourses using Algorithm 1, and then project the recourse to the feasible region to satisfy the constraints posed by categorical features.
@@ C.4 Actionable Features and Post-processing
C.5 Baselines for Smoothness Results
In this section, we revisit the smoothness results in Section 4.1 by including a baseline. For any $\beta$, the optimization problem in Equation 7 can be solved by the type of gradient-based techniques that are common in adversarial training [Madry et al., 2018]. Since ROAR is one such algorithm, we can modify it by changing its objective function and use it as a baseline to Algorithm 1.
@@ C.5 Baselines for Smoothness Results
C.6 Computing Robust Recourse
@@
In this section, we repeat the experiments in Section 4.2 but also include comparisons with RBR [Nguyen et al., 2022]. Figure 10 depicts the trade-off between worst-case validity and cost for all datasets (rows) and models (columns). Since RBR does not have the same parameters as our algorithm and ROAR, the trade-off for RBR is obtained by replicating their experiments and setting the ambiguity sizes to $\epsilon_{0},\epsilon_{1}\in[0,1]$ with increments of 0.5, and the maximum recourse cost $\delta=\|x_{0}-x_{r}\|_{1}+\delta_{+}$ to $\delta_{+}\in[0,1]$ with increments of 0.2.
For neural network models, the fidelity of LIME approximation for Synthetic, Small Business, and German datasets is 0.93, 0.54, and 0.02 when averaged over instances where recourse is provided. These fidelity values represent a wide range for the approximation quality of LIME. We see, in the right panel of Figure 10, that our algorithm performs similarly to or better than RBR when fidelity is relatively reasonable, and only when fidelity is extremely low (German dataset) does its performance become dominated by RBR. For logistic regression models, in the left panel of Figure 10, our algorithm generally finds very high validity, while this is not the case for RBR, which sometimes achieves validity even smaller than ROAR. We also observe that, to achieve lower or moderate validity values, $\lambda$ needs to be set higher than 0.3 for our algorithm.
C.7 Experiments on Larger Datasets
@@
We use the ACSIncome-CA [Ding et al., 2021] dataset for experiments in this section. This dataset originally consisted of 195,665 data points and 10 features, 7 of which are categorical and have been one-hot encoded. However, to lower the runtime, we subsampled the dataset to include 50,000 data points and removed the categorical feature “occupation (OCCP), as it contains more than 500 different occupations. This resulted in more than 250 features after one-hot encoding.